OpenStudy (anonymous):

(s+3)+[(7s-2)+3]=52 s=6 I got that is was true is this so?

OpenStudy (anonymous):

yes s=6 is correct.

OpenStudy (anonymous):

\[s+3+[(7s-2)+3]=52\] \[s+3+7s+1=52\] \[8s+4=52\] \[8s=52-4\] \[8s=48\] \[s=6\]